# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ 2.B. Distributional LS-category
denote the space of probability measures on $Z$ supported by at most $n$ points, equipped with the Lévy–Prokhorov metric [Pr] (or the Wasserstein distance metric), see also [DJ, Section 3.1]. Given the pointed metric space $(Z,z_{0})$, for each $z\in Z$, let $P(z,z_{0})=\{\phi\in P(Z)\mid\phi(0)=z\}$.
@@
The distributional LS-category of a map $f:(X,x_{0})\to(Y,y_{0})$, denoted $\operatorname{\mathsf{dcat}}(f)$, is the smallest integer $n$ for which there exists a continuous map $H:X\to{\mathcal{B}}_{n+1}(P(Y))$ such that $H(x)\in{\mathcal{B}}_{n+1}(P(f(x),y_{0}))$ for all $x\in X$.
Upon taking $f$ as the identity map, we get the notion of the distributional LS-category of a space $X$, denoted $\operatorname{\mathsf{dcat}}(X)$. For the following discussion, we refer to [DJ, Sections 3.2 and 4.2].
Given a space $X$, $\operatorname{\mathsf{dcat}}(X)$ is a lower bound to its classical LS-category $\operatorname{\mathsf{cat}}(X)$. It is easy to see that $\operatorname{\mathsf{dcat}}(X)=0$ if and only if $X$ is contractible. Hence, $\operatorname{\mathsf{dcat}}(S^{n})=1$ for all $n$. Interestingly, $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{n})=1$ for all $n$, in contrast to the well-known equality $\operatorname{\mathsf{cat}}({\mathbb{R}}P^{n})=n$. This result leads to various differences between the theories of $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$, see [DJ, Section 6.1] and [J2, Remark 4.13]. In general, $\operatorname{\mathsf{dcat}}(X)$ can be arbitrarily large; indeed, the rational and the real cup lengths of $X$ are lower bounds to $\operatorname{\mathsf{dcat}}(X)$. However, we note that, unlike the case of the classical LS-category, the $R$-cup length of $X$, denoted $c\ell_{R}(X)$, is not necessarily a lower bound to $\operatorname{\mathsf{dcat}}(X)$ for a general ring $R$ due to the example of ${\mathbb{R}}P^{n}$. Furthermore, $\operatorname{\mathsf{dcat}}$ is a homotopy invariant of spaces. So, for a finitely generated discrete group $\Gamma$, one can define $\operatorname{\mathsf{dcat}}(\Gamma):=\operatorname{\mathsf{dcat}}(B\Gamma)$, where $B\Gamma$ denotes the unique (up to homotopy) classifying space of $\Gamma$.
The well-known result of Eilenberg and Ganea [EG] says that $\operatorname{\mathsf{cat}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma)$, where $\operatorname{{\rm cd}}(\Gamma)=\max\{k\in\mathbb{Z}\mid H^{k}(\Gamma,{\mathcal{A}})\neq 0\text{ for some }{\mathbb{Z}}\Gamma\text{-module }{\mathcal{A}}\}$ is the cohomological dimension of $\Gamma$. For any finite group $\Gamma$, while $\operatorname{\mathsf{cat}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma)$ is infinite, it follows from [KW, Theorem 7.2] that $\operatorname{\mathsf{dcat}}(\Gamma)\leq|\Gamma|-1$ (see also [J1, Section 8.A]). In fact, for any prime $p\geq 2$, $\operatorname{\mathsf{dcat}}({\mathbb{Z}}_{p})=p-1$ due to [Dr2, Theorem 5.12] (see also [DJ] for the case $p=2$). Therefore, an analog of the Eilenberg–Ganea theorem does not hold in the theory of distributional LS-category for groups having torsion. However, such an analog holds for torsion-free discrete groups.
@@
We note that the distributional LS-category of a space is a special case of the general notion of the “distributional sectional category” of a fibration, denoted $\mathsf{dsecat}$, which is the distributional analog of the classical notion of the sectional category [Sch][Ja]. We refer the reader to [J1, Section 5] for details on $\mathsf{dsecat}$.
@@
Here, $\delta_{n}:Y\to SP^{n!}(Y)$ is the diagonal embedding.
@@
The following covering map inequality for $\operatorname{\mathsf{dcat}}$ (which is a direct analog of the corresponding inequality for $\operatorname{\mathsf{cat}}$, see [CLOT, Corollary 1.45]) will be very useful in Sections 5.A and 8.A.
@@ 2.C. Topological basics
Given a fibration $f:E\to B$ where $B$ is a CW complex and the fiber $F$ is $(n-2)$-connected, the primary obstruction to a section of $f$ is a cohomology class $\kappa_{n}\in H^{n}(B;H_{n-1}(F))$, where coefficients are in the $\pi_{1}(B)$-module $H_{n-1}(F)$.
For a discrete group $\Gamma$, it follows from [Sch] that the $n^{th}$ power of its Berstein–Schwarz class, $\beta_{\Gamma}^{n}$, is the primary obstruction to a section of the $n^{th}$ Ganea fibration $p_{\Gamma}:G_{n}(B\Gamma)\to B\Gamma$ defined in Section 2.A.
@@
We end this section by recalling Poincaré duality. Given a closed manifold $M$ of dimension $n$, the orientation sheaf ${\mathcal{O}}_{M}$ on $M$ defines a fundamental class $[M]$ of $M$ such that for any local coefficients ${\mathcal{A}}$ on $M$ and integer $k\in\{0,\ldots,n\}$, the $k^{th}$ homomorphism
$\text{PD}_{k}:H^{k}(M;{\mathcal{A}})\to H_{n-k}(M;{\mathcal{A}}\otimes{\mathcal{O}}_{M})$
that sends each $\alpha\in H^{k}(M;{\mathcal{A}})$ to the cap product $[M]\frown\alpha\in H_{n-k}(M;{\mathcal{A}}\otimes{\mathcal{O}}_{M})$ is an isomorphism, see [Bre, Chapter V, Sections 9 and 10]. If a mapping $f:M\to N$ takes the orientation sheaf ${\mathcal{O}}_{N}$ to ${\mathcal{O}}_{M}$, then the degree of $f$, denoted $\text{deg}(f)$, is the integer for which $f_{*}([M])=\text{deg}(f)[N]$. If $\text{deg}(f)=\pm 1$, then $f$ induces an epimorphism $f_{\bullet}:\pi_{1}(M)\to\pi_{1}(N)$ of the fundamental groups.
@@ 3. For essential manifolds
    There exists a coefficient system ${\mathcal{A}}$ such that the induced homomorphism $f^{*}:H^{n}(B\pi;{\mathcal{A}})\to H^{n}(M;f^{*}({\mathcal{A}}))$ is non-trivial.
@@
So, it is natural to look for essential manifolds whose distributional LS-category is equal to the dimension of the manifold.
3.A. Torsion-free fundamental groups
@@
Let us formalize the cap property. The orientation sheaf ${\mathcal{O}}_{M}$ on $M$ is uniquely determined by the kernel of the homomorphism $w:\pi\to{\mathbb{Z}}_{2}=\text{Aut}({\mathbb{Z}})$, where $w$ is the first Stiefel–Whitney class, see [Ba, Page 32]. Hence, for $M$ to take (co)homology with coefficients in ${\mathcal{O}}_{M}$ is the same as taking (co)homology with coefficients in the twisted integers ${\mathbb{Z}}_{w}$. Since $[M]\in H_{n}(M;{\mathcal{O}}_{M})=H_{n}(M;{\mathbb{Z}}_{w})$ and $f:M\to B\pi$ induces an isomorphism of the fundamental groups, we have that $f_{*}([M])\in H_{n}(B\pi;{\mathbb{Z}}_{w})$. Therefore, $f_{*}([M])\frown\kappa_{n}\in H_{0}(B\pi;{\mathbb{Z}}_{w}\otimes{\mathcal{F}}).$
3.4 Proposition.
Let $M$ be a closed essential $n$-manifold with torsion-free fundamental group $\pi$. If $M$ satisfies the cap property, then $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)=n$.
Proof.
@@
due to Poincaré duality in $B\pi$ in ${\mathcal{O}}_{\pi}\otimes{\mathcal{F}}$ coefficients. Thus, $M$ satisfies the cap property. Hence, the conclusion follows from Proposition 3.4. ∎
@@
3.B. Symplectically aspherical manifolds
To emphasize the utility of Corollary 3.5, we will present in this section various examples of non-aspherical essential manifolds whose distributional LS-category turns out to be maximum.
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$ such that $[\omega]^{n}\neq 0$, where $[\omega]\in H^{2}(M;{\mathbb{R}})$ is the de Rham cohomology class corresponding to the non-degenerate symplectic $2$-form $\omega$. If $(M,\omega)$ is simply connected, then clearly, $n\leq c\ell_{{\mathbb{R}}}(M)\leq\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)\leq n$ implies $\operatorname{\mathsf{dcat}}(M)=n$. So, we assume that $M$ is not simply connected. Let $c_{1}\in H^{2}(M;{\mathbb{Z}})$ denote the first Chern class of $M$ and let $h:\pi_{2}(M)\to H_{2}(M;{\mathbb{Z}})$ denote the Hurewicz homomorphism. We follow [Ori] and call $(M,\omega)$ spherically monotone if there exists some $\lambda\in{\mathbb{R}}$ such that the composition $([\omega]-\lambda c_{1})\circ h:\pi_{2}(M)\to H_{2}(M;{\mathbb{Z}})\to{\mathbb{R}}$ is trivial. When $\lambda=0$, then $(M,\omega)$ is called symplectically aspherical. Examples of these kinds include symplectic manifolds $M$ for which $\pi_{2}(M)=0$, such as aspherical manifolds.
@@ 3.B. Symplectically aspherical manifolds
We note that the LS-category of these symplectic manifolds was calculated in [RO] and [Ori] using the notion of category weight (see [R1]). For $\operatorname{\mathsf{dcat}}$, the author has not been able to develop a “useful” notion of “distributional category weight” that can determine $\operatorname{\mathsf{dcat}}(M)$. However, our current technique does give many new computations for $\operatorname{\mathsf{dcat}}$ of symplectically aspherical manifolds.
@@
However, to get new computations of $\operatorname{\mathsf{dcat}}$ using Corollary 3.7, we are interested in some non-aspherical SA manifolds (while Gompf [Go] provides such examples in dimension $4$, we don’t know if they satisfy the cap property and so, their $\operatorname{\mathsf{dcat}}$ cannot be determined by our methods). Indeed, many such manifolds exist in all even dimensions $\geq 4$ and we obtain new computations of $\operatorname{\mathsf{dcat}}$ in all those cases.
@@
for some $h\geq g$ with $\pi_{1}(X)=\Gamma_{g}$ and $\Omega\in H^{2}(X;{\mathbb{R}})$ such that $\langle\Omega,\phi_{*}[S^{2}]\rangle=0$ and $\psi_{y}^{*}(\Omega)\neq 0$ for all maps $\phi:S^{2}\to X$ and $\psi_{y}:f^{-1}(y)\hookrightarrow X$ for $y\in S^{2}$. For any $k\geq 1$, let $Y_{k}=\Sigma_{h}\times\Sigma_{k}$. Then the Gompf symplectic fiber sum $X\#_{\Sigma_{h}}Y_{k}$ (see [IKRT, Section 3] for its definition and also Section 6.A for a general description of fiber sums) is SA by [KRT, Proposition 3.3]. Due to [KRT, Corollary 4.5], $\pi_{1}(X\#_{\Sigma_{h}}Y_{k})=\Gamma_{g}\oplus\Gamma_{k}$. Hence, we apply Corollary 3.7 to get $\operatorname{\mathsf{dcat}}(X\#_{\Sigma_{h}}Y_{k})=4$. We note that in general, the fiber sums $X\#_{\Sigma_{h}}Y_{k}$ need not be aspherical.
@@
For any $n\geq 2$, let $\Gamma^{n}:=\oplus_{i=1}^{n}\Gamma_{g_{i}}$, where $\Gamma_{g_{i}}=\pi_{1}(\Sigma_{g_{i}})$ for $g_{i}\geq 1$. We saw in Examples 3.10 and 3.12 that if at least two summands in $\Gamma^{n}$ are ${\mathbb{Z}}^{2}$, then there exists a non-aspherical SA $2n$-manifold $M$ such that $\pi_{1}(M)=\Gamma^{n}$ and $\operatorname{\mathsf{dcat}}(M)=2n$. Now, let us assume that at most one summand in $\Gamma^{n}$ is ${\mathbb{Z}}^{2}$. Then there exists $j\leq n$ such that $g_{j}\geq 2$. In Example 3.13, we take $g=g_{j}$ and $k=g_{s}$ for any fixed $s\neq j$ and get $M_{js}=X\#_{\Sigma_{h}}Y_{g_{s}}$ for some $h\geq g_{j}$. If $n\geq 3$, then we let ${\mathcal{P}}_{n-2}=\prod_{i=1,i\neq j,s}^{n}\Sigma_{g_{i}}$ and consider $M=M_{js}\times{\mathcal{P}}_{n-2}$. Whenever $M_{js}\neq\Sigma_{g_{j}}\times\Sigma_{g_{s}}$, it is not aspherical. So, in those cases, $M$ is a non-aspherical SA $2n$-manifold such that $\pi_{1}(M)=\Gamma^{n}$ and $\operatorname{\mathsf{dcat}}(M)=2n$ due to Corollary 3.7.

### v3

@@ 2.2. Distributional category
denote the space of probability measures on $Z$ supported by at most $n$ points, equipped with the Lévy–Prokhorov metric [Pr] (or the Wasserstein distance metric), see also [DJ, Section 3.1]. Here, $\operatorname{{\rm supp}}(\mu)$ denotes the support of $\mu$. Given the pointed metric space $(Z,z_{0})$, for each $z\in Z$, let $P(z,z_{0})=\{\phi\in P(Z)\mid\phi(0)=z\}$.
@@
The distributional category of a map $f:(X,x_{0})\to(Y,y_{0})$, denoted $\operatorname{\mathsf{dcat}}(f)$, is the smallest integer $n$ for which there exists a continuous map $H:X\to{\mathcal{B}}_{n+1}(P(Y))$ such that $H(x)\in{\mathcal{B}}_{n+1}(P(f(x),y_{0}))$ for all $x\in X$.
Upon taking $f$ as the identity map, we get the notion of the distributional category of a space $X$, denoted $\operatorname{\mathsf{dcat}}(X)$. For the following discussion, we refer to [DJ, Sections 3.2 and 4.2].
Given a space $X$, $\operatorname{\mathsf{dcat}}(X)$ is a lower bound to its classical LS-category $\operatorname{\mathsf{cat}}(X)$. It is easy to see that $\operatorname{\mathsf{dcat}}(X)=0$ if and only if $X$ is contractible. Hence, $\operatorname{\mathsf{dcat}}(S^{n})=1$ for all $n$. Interestingly, $\operatorname{\mathsf{dcat}}({\mathbb{R}}P^{n})=1$ for all $n$, in contrast to the well-known equality $\operatorname{\mathsf{cat}}({\mathbb{R}}P^{n})=n$. This result leads to various differences between the theories of $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$, see [DJ, Section 6.1] and [J2, Remark 4.13]. In general, $\operatorname{\mathsf{dcat}}(X)$ can be arbitrarily large; indeed, the rational and the real cup-lengths of $X$ are lower bounds to $\operatorname{\mathsf{dcat}}(X)$. However, we note that, unlike the case of the classical LS-category, the $R$-cup-length of $X$, denoted $c\ell_{R}(X)$, is not necessarily a lower bound to $\operatorname{\mathsf{dcat}}(X)$ for a general ring $R$ due to the example of ${\mathbb{R}}P^{n}$. Furthermore, $\operatorname{\mathsf{dcat}}$ is a homotopy invariant of spaces. So, for a finitely generated discrete group $\Gamma$, one can define $\operatorname{\mathsf{dcat}}(\Gamma):=\operatorname{\mathsf{dcat}}(B\Gamma)$, where $B\Gamma$ denotes the unique (up to homotopy) classifying space of $\Gamma$.
The well-known result of Eilenberg and Ganea [EG] says that $\operatorname{\mathsf{cat}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma)$, where $\operatorname{{\rm cd}}(\Gamma)=\max\{k\in\mathbb{Z}\mid H^{k}(\Gamma,{\mathcal{A}})\neq 0\text{ for some }{\mathbb{Z}}\Gamma\text{-module }{\mathcal{A}}\}$ is the cohomological dimension of $\Gamma$. For any finite group $\Gamma$, while $\operatorname{\mathsf{cat}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma)$ is infinite, it follows from [KW, Theorem 7.2] that $\operatorname{\mathsf{dcat}}(\Gamma)\leq|\Gamma|-1$ (see also [J1, Section 8.A]). In fact, for any prime $p\geq 2$, $\operatorname{\mathsf{dcat}}({\mathbb{Z}}_{p})=p-1$ due to [Dr2, Theorem 5.12] (see also [DJ] for the case $p=2$). Therefore, an analog of the Eilenberg–Ganea theorem does not hold in the theory of distributional category for groups having torsion. However, such an analog holds for torsion-free discrete groups.
@@
We note that the distributional category of a space is a special case of the general notion of the “distributional sectional category” of a fibration, denoted $\mathsf{dsecat}$, which is the distributional analog of the classical notion of the sectional category [Sch][Ja]. We refer the reader to [J1, Section 5] for details on $\mathsf{dsecat}$.
@@
Here, $\delta_{n}:Y\to SP^{n!}(Y)$ is the diagonal embedding into the $n!$ symmetric power of $Y$, which is defined as the orbit space of the natural action of the symmetric group $S_{n!}$ on the product space $Y^{n!}$.
@@
The following covering map inequality for $\operatorname{\mathsf{dcat}}$ (which is a direct analog of the corresponding inequality for $\operatorname{\mathsf{cat}}$, see [CLOT, Corollary 1.45]) will be very useful in Sections 5.1 and 8.1.
@@ 2.3. Topological basics
Given a fibration $f:E\to B$, where $B$ is a CW complex and the fiber $F$ is $(n-2)$-connected for some $n\geq 3$, the primary obstruction to a section of $f$ is a cohomology class $\kappa_{n}\in H^{n}(B;H_{n-1}(F))$, where coefficients are in the $\pi_{1}(B)$-module $H_{n-1}(F)$.
For a discrete group $\Gamma$, it follows from [Sch] that the $n$-th power of its Berstein–Schwarz class, $\beta_{\Gamma}^{n}$, is the primary obstruction to a section of the $n$-th Ganea fibration $p_{\Gamma}:G_{n}(B\Gamma)\to B\Gamma$ defined in Section 2.1.
@@
We end this section by recalling Poincaré duality. Given a closed manifold $M$ of dimension $n$, the orientation sheaf ${\mathcal{O}}_{M}$ on $M$ defines a fundamental class $[M]$ of $M$ such that for any local coefficients ${\mathcal{A}}$ on $M$ and integer $k\in\{0,\ldots,n\}$, the $k$-th homomorphism
$\text{PD}_{k}:H^{k}(M;{\mathcal{A}})\to H_{n-k}(M;{\mathcal{A}}\otimes{\mathcal{O}}_{M}),$
that sends each $\alpha\in H^{k}(M;{\mathcal{A}})$ to the cap product $[M]\frown\alpha\in H_{n-k}(M;{\mathcal{A}}\otimes{\mathcal{O}}_{M})$, is an isomorphism, see [Bre, Chapter V, Sections 9 and 10]. If a mapping $f:M\to N$ takes the orientation sheaf ${\mathcal{O}}_{N}$ to ${\mathcal{O}}_{M}$, then the degree of $f$, denoted $\text{deg}(f)$, is the integer for which $f_{*}([M])=\text{deg}(f)[N]$. If $\text{deg}(f)=\pm 1$, then $f$ induces an epimorphism $f_{\bullet}:\pi_{1}(M)\to\pi_{1}(N)$ of the fundamental groups.
@@ 3. For essential manifolds
    There exists a coefficient system ${\mathcal{A}}$ and a map $f:M\to B\pi$ such that the induced homomorphism $f^{*}:H^{n}(B\pi;{\mathcal{A}})\to H^{n}(M;f^{*}({\mathcal{A}}))$ is non-trivial.
If ${\mathcal{A}}={\mathbb{Q}}$ in part (3) above, then $M$ is called rationally essential.
@@
So, it is natural to look for essential manifolds whose distributional category is equal to the dimension of the manifold.
3.1. Torsion-free fundamental groups
@@
Let us formalize the cap property. The orientation sheaf ${\mathcal{O}}_{M}$ on $M$ is uniquely determined by the kernel of the homomorphism $w:\pi\to{\mathbb{Z}}_{2}=\text{Aut}({\mathbb{Z}})$, where $w$ is the first Stiefel–Whitney class of $M$, see [Ba, Page 32]. Hence, for $M$ to have (co)homology with coefficients in ${\mathcal{O}}_{M}$ is the same as having (co)homology with coefficients in the twisted integers ${\mathbb{Z}}_{w}$. Since $[M]\in H_{n}(M;{\mathcal{O}}_{M})=H_{n}(M;{\mathbb{Z}}_{w})$ and $f:M\to B\pi$ induces an isomorphism of the fundamental groups, we have that $f_{*}([M])\in H_{n}(B\pi;{\mathbb{Z}}_{w})$. Therefore, $f_{*}([M])\frown\kappa_{n}\in H_{0}(B\pi;{\mathbb{Z}}_{w}\otimes{\mathcal{F}})$, where ${\mathbb{Z}}_{w}\otimes{\mathcal{F}}$ is a ${\mathbb{Z}}\pi$-module.
Proof of Theorem A.
@@
due to Poincaré duality in $B\pi$ in ${\mathcal{O}}_{\pi}\otimes{\mathcal{F}}$ coefficients. Thus, $M$ satisfies the cap property. Hence, the conclusion follows from Theorem A. ∎
@@
3.2. Symplectically aspherical manifolds
To emphasize the utility of Corollary 3.4, we will present in this section various examples of non-aspherical essential manifolds whose distributional category turns out to be maximum.
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$ such that $[\omega]^{n}\neq 0$, where $[\omega]\in H^{2}(M;{\mathbb{R}})$ is the de Rham cohomology class corresponding to the non-degenerate symplectic $2$-form $\omega$. If $(M,\omega)$ is simply connected, then clearly, $n\leq c\ell_{{\mathbb{R}}}(M)\leq\operatorname{\mathsf{dcat}}(M)\leq\operatorname{\mathsf{cat}}(M)\leq n$ implies $\operatorname{\mathsf{dcat}}(M)=n$. So, we assume that $M$ is not simply connected. Let $c_{1}\in H^{2}(M;{\mathbb{Z}})$ denote the first Chern class of $M$ and let $h:\pi_{2}(M)\to H_{2}(M;{\mathbb{Z}})$ denote the Hurewicz homomorphism. We follow [Ori] and call $(M,\omega)$ spherically monotone if there exists some $\lambda\in{\mathbb{R}}$ such that the composition
$([\omega]-\lambda c_{1})\circ h:\pi_{2}(M)\to H_{2}(M;{\mathbb{Z}})\to{\mathbb{R}}$
is trivial. When $\lambda=0$, then $(M,\omega)$ is called symplectically aspherical. Examples of these kinds include symplectic manifolds whose second homotopy groups vanish, such as aspherical manifolds.
@@ 3.2. Symplectically aspherical manifolds
We note that the LS-category of these symplectic manifolds was calculated in [RO] and [Ori] using the notion of category weight (see [St][R1]). For $\operatorname{\mathsf{dcat}}$, the author has not been able to develop a “useful” notion of “distributional category weight” that can determine $\operatorname{\mathsf{dcat}}(M)$. However, our current technique does give many new computations for $\operatorname{\mathsf{dcat}}$ of symplectically aspherical manifolds.
@@
However, to get new computations of $\operatorname{\mathsf{dcat}}$ using Corollary 3.6, we are interested in non-aspherical SA manifolds (while Gompf [Go] provides such examples in dimension $4$, we don’t know if they satisfy the cap property and so, their $\operatorname{\mathsf{dcat}}$ cannot be determined by our methods). Indeed, many such manifolds exist in all even dimensions $\geq 4$ and we obtain new computations of $\operatorname{\mathsf{dcat}}$ in all those cases.
@@
for some $h\geq g$ with $\pi_{1}(X)=\Gamma_{g}$ and $\Omega\in H^{2}(X;{\mathbb{R}})$ such that $\langle\Omega,\phi_{*}[S^{2}]\rangle=0$ and $\psi_{y}^{*}(\Omega)\neq 0$ for all smooth maps $\phi:S^{2}\to X$ and $\psi_{y}:f^{-1}(y)\hookrightarrow X$ for $y\in S^{2}$. For any $k\geq 1$, let $Y_{k}=\Sigma_{h}\times\Sigma_{k}$. Then the Gompf symplectic fiber sum $X\#_{\Sigma_{h}}Y_{k}$ (see [IKRT, Section 3] for its definition and also Section 6.1 for a general description of fiber sums) is SA by [KRT, Proposition 3.3]. Due to [KRT, Corollary 4.5],
$\pi_{1}\left(X\#_{\Sigma_{h}}Y_{k}\right)=\Gamma_{g}\oplus\Gamma_{k}.$
Hence, we apply Corollary 3.6 to get $\operatorname{\mathsf{dcat}}(X\#_{\Sigma_{h}}Y_{k})=4$. We note that in general, the fiber sums $X\#_{\Sigma_{h}}Y_{k}$ need not be aspherical.
We now explain how for any $n\geq 2$, any direct sum of $n$-number of surface groups (i.e., fundamental groups of closed orientable surfaces $\Sigma_{g}$ of genus $g\geq 1$) can be realized as the fundamental group of a closed symplectically aspherical $2n$-manifold $M$ that is not aspherical but whose $\operatorname{\mathsf{dcat}}(M)=\operatorname{\mathsf{cat}}(M)=2n$.
@@
For any $n\geq 2$, let $\Gamma^{n}:=\oplus_{i=1}^{n}\Gamma_{g_{i}}$, where $\Gamma_{g_{i}}=\pi_{1}(\Sigma_{g_{i}})$ for $g_{i}\geq 1$. We saw in Examples 3.9 and 3.11 that if at least two summands in $\Gamma^{n}$ are ${\mathbb{Z}}^{2}$, then there exists a non-aspherical SA $2n$-manifold $M$ such that $\pi_{1}(M)=\Gamma^{n}$ and $\operatorname{\mathsf{dcat}}(M)=2n$. Now, let us assume that at most one summand in $\Gamma^{n}$ is ${\mathbb{Z}}^{2}$. Then there exists $j\leq n$ such that $g_{j}\geq 2$. In Example 3.12, we take $g=g_{j}$ and $k=g_{s}$ for any fixed $s\neq j$ and get $M_{js}=X\#_{\Sigma_{h}}Y_{g_{s}}$ for some $h\geq g_{j}$. If $n\geq 3$, then we let ${\mathcal{P}}_{n-2}=\prod_{i=1,i\neq j,s}^{n}\Sigma_{g_{i}}$ and consider $M=M_{js}\times{\mathcal{P}}_{n-2}$. Whenever $M_{js}\neq\Sigma_{g_{j}}\times\Sigma_{g_{s}}$, it is not aspherical. So, in those cases, $M$ is a non-aspherical SA $2n$-manifold such that $\pi_{1}(M)=\Gamma^{n}$ and $\operatorname{\mathsf{dcat}}(M)=2n$ due to Corollary 3.6.